Solve this system of equations using substitution. Put your answer in ordered pair for, (x, y). x + 2y = 7 2x + y = 8
Answer by mananth(16946)

2x+y=8 .............1
y=8-2x
x+2y=7
Substitute y in (2)
x+2(8-2x)=7
x+16-4x =7
x-4x=-9
-3x=-9
/-3
x=3
Plug the value of x in (1)
2x+y=8
2*3+y= 8
6+y=8
y=2
/1
y=2
The solution is (3,2)
m.ananth@hotmail.ca